OpenStudy (anonymous):

Integration by parts: the integral of ln(2x + 1) dx I am stuck after trying the formula. please help me.

OpenStudy (anonymous):

Let Int denote the integral sign. Integration by parts is Int[u(dv)] = u*v - Int[v(du)] Let u =ln( 2x + 1) so du = [2(dx)]/[2x+1] dv = dx so v = x uv = x*ln(2x+1) Int[v(du)] = Int[(2x)/(2x+1)](dx) Since 2x/(2x+1) = 1 - 1/(2x+1) we split it up into Int[(dx)] - Int[(1/(2x+1))(dx)] Int[(dx)] = x Int[(1/(2x+1))(dx)] = (ln(2x+1))/2 The integral is then x*ln(2x+1) - x + ln(2x+1)/2 Which can be verified by taking the derivative

OpenStudy (anonymous):

You want to start off with a substitution. Let p=2x+1 then dp=2dx Rewrite the integral this way: \[\frac{1}{2}\int\limits \ln(p)dp\] From here, let u=ln(p) and dv=dp so du=1/p and v=p Using the formula you have: \[\frac{1}{2}(p \ln(p)-\int\limits \frac{p}{p}dx)\] This reduces to:\[\frac{1}{2}(p \ln(p)-\int\limits dp)=\frac{1}{2}(p \ln(p)-p)\] Now replug in your p: \[\frac{1}{2}(2x+1)\ln(2x+1)-\frac{1}{2}(2x+1)+C\] Does that make sense?

OpenStudy (anonymous):

above answer is absolutely right. but if i were you (taking a calc course) i would just try to remember that \[\int\ln(x)dx=x\ln(x)-x\] then if you know that the above is a simple u-sub. integrating the natural log is a question math teachers love to ask (parts is the way) so i would just memorize it.